### v2

Abstract
We introduce and analyze a mesh-free two-level hybrid Tucker tensor format for approximating multivariate functions, which combines the product Chebyshev interpolation with the alternating least-squares (ALS) based Tucker decomposition of the tensor of Chebyshev coefficients. This construction allows to avoid the expensive rank-structured grid-based approximation of function-related tensors on large spatial grids, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients. Thus, we can compute the nearly optimal Tucker decomposition of the 3D function with controllable accuracy $\varepsilon>0$ without discretizing the function on the full grid in the domain, but only using its values at small set of Chebyshev nodes. Finally, we can represent the function in the algebraic Tucker format with optimal $\varepsilon$-rank on an arbitrarily large 3D tensor grid in the computational domain by discretizing the Chebyshev polynomials on that grid. The rank parameters of the Tucker-ALS decomposition of the coefficient tensor can be much smaller than the polynomial degrees of the initial Chebyshev interpolant obtained via a function independent polynomial basis set. It is shown that our techniques can be gainfully applied to the long-range part of the singular electrostatic potential of multi-particle systems approximated in the range-separated tensor format. We provide error and complexity estimates and demonstrate the computational efficiency of the proposed techniques on challenging examples, including the multi-particle electrostatic potential for large bio-molecular systems and lattice-type compounds.
keywords:
Tensor numerical methods, Tucker and canonical tensor formats, Chebyshev polynomial interpolation, multilinear algebra, multi-particle electrostatic potentials.
{MSCcodes}
65F30, 65F50, 65N35, 65F10
1 Introduction
Tensor numerical methods (TNM) intrinsically simplify the numerical solution of multi-dimensional problems to one-dimensional calculations. In recent years, TNM are proven to be a powerful tool for solving numerically intensive multi-dimensional problems in scientific computing arising in computational quantum chemistry [34], multi-dimensional steady state and dynamical PDE-driven problems, data science, stochastic simulations, optimal control problems, etc., see, e.g., [1, 37] for a detailed discussion and related references. The benefits of TNM for $d$-dimensional problems are due to the reduction of both computational and storage complexity to merely linear scaling in $d$ based on the representation of functions and operators in low-parametric rank structured tensor formats.
TNM originated as bridging of approximation theory for multi-dimensional functions and operators from one side and the modern multilinear algebra techniques from the other side. The construction of efficient TNM is guided by the trade-off between approximation accuracy and computational complexity of the chosen rank-structured approximation algorithms, while taking into account the compatibility of the resultant tensor representations with available multilinear algebra. The total numerical complexity of tensor approximations is a sum of the high computational (and storage) cost for (a) the algebraic operations needed for the low-rank approximation of multi-dimensional functions and operators in the conventional canonical (CP) [30], Tucker [60], and rather recent tensor train (TT) [48], quantized TT (QTT) [36] or range-separated (RS) [6] tensor formats, and (b) the cost of function evaluations required for implementation of the tensor algorithms. Thereafter, the overall approximation accuracy can be controlled by rank parameters as well as the precision of the chosen functional approximations (given the storage budget, accuracy is limited by the data regularity). A number of recent developments confirm efficiency of TNM in different research fields [2, 42, 55, 44, 43, 52, 51, 61].
TNM are usually applied to grid-based rank-structured tensor approximation of target functions and operators discretized on large $n^{\otimes d}$ spatial grids in $\mathbb{R}^{d}$, $d\geq 2$, for the numerical solution of large scale scientific computing problems [37, 34, 7]. In this direction the sparse grid methods, stochastic homogenization techniques and UQ approaches also provide powerful computational tools for special classes of problems. Alternatively, classical mesh-free approximation methods for functions are based on their representation in problem-independent basis sets, such as trigonometric functions (Fourier), (piecewise) polynomials, sinc functions or Gaussians [53, 59, 13, 47, 44, 24]. The fast multipole method provides another powerful example of a mesh-free approach for certain kernel-based problems [25]. The other well-developed computational technique is based on the general principle of model reduction [4, 8].
This paper is partially motivated by recent advances in tensor techniques for 3D Chebyshev interpolation [3, 20, 22, 29, 54]. Chebyshev polynomials [16] are among the most commonly used set of interpolating functions due to many beneficial features [10, 9, 47]. Error analysis for tensor product interpolation by Chebyshev polynomials was presented for example in [27]. MATLAB implementations of the Chebyshev interpolation method up to 3D are available within the chebfun package [59, 22, 29, 57, 58, 56, 12]. The basic version of the trivariate chebfun3 [29] applies to ${\bf T}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ represented in Tucker format, which contains the values of the input function at Chebyshev nodes, thus using only chebfun for 1D polynomial interpolation. Approximation of the target tensor $\bf T$} in Tucker format is performed via a variant of the heuristic method adaptive cross approximation (ACA) [3] that combines ACA for the 2D slices with agglomeration of slices along the third spatial direction. Hence, the number of functional calls in this approach is $\mathcal{O}(mr^{2})$ and may be beneficial for low-rank smooth functions. Here $m$ is the maximum Chebyshev degrees in the three dimensions, and $r$ is the Tucker rank. In contrast, the direct tensor product Chebyshev interpolant requires $m^{3}$ function calls. A modification of chebfun3 that reduces the number of function calls was presented in [20].
While powerful, these interpolatory methods face a significant challenge. For functions with sufficient smoothness, this approach can achieve quasi-optimal accuracy with a relatively small basis set. However, for functions with low regularity or localized features, such as those with sharp gradients or singularities, achieving a satisfactory accuracy may require a very large number of interpolating functions (i.e., a high polynomial degree $m$), which makes further use of such interpolants in scientific computing non-tractable. That is the main bottleneck for approximation of multivariate functions by Chebyshev polynomials in large scale scientific computing involving low regularity functions.
In this paper, we introduce and analyze a novel method that directly addresses this challenge, which combines the tensor product Chebyshev polynomial interpolation in the computational domain with the ALS-based Tucker decomposition of the rather small tensor of Chebyshev coefficients [^1] Using non-orthogonal interpolating functions leads to small modification of the computational scheme., termed Chebyshev-Tucker (ChebTuck) format. This leads to the construction of mesh-free two-level Tucker tensor format for approximation of multivariate functions, either given explicitly in analytic form or only available in the form of discrete grid-based representations. It allows to avoid the expenses of the rank-structured approximation of function-related tensors defined on fine spatial grids in $\mathbb{R}^{3}$, while benefiting from the Tucker decomposition of the rather small core tensor of Chebyshev coefficients, which leads to nearly optimal rank parameters as for the well established grid-based methods. The $\varepsilon$-ranks inherited from the Tucker-ALS decomposition are almost optimal and can be much smaller than the polynomial degree, $m$, of the initial Chebyshev interpolant which, in turn, is chosen independent of the grid sizes in grid-based methods. Thus, we compute the nearly optimal Tucker decomposition of the 3D function without discretizing the function on the full grid in the computational domain, but only using its values at Chebyshev nodes. Note that in the case of functions with multiple cusps (or even stronger singularities), our method applies to the long-range component of the target function approximated in the range-separated (RS) tensor format [6].
We underline that, given accuracy $\varepsilon>0$, our main goal in this paper is to construct numerical methods that allow to compute the mesh-free Tucker tensor approximation of the target function with almost optimal tensor rank ${\bf r}=(r_{1},r_{2},r_{3})$ which is usually much smaller than the Chebyshev polynomials degree $m$ that guarantees the accuracy $\varepsilon$ of the Chebyshev interpolant. In turn, the parameter $m$ is supposed to be much smaller than the grid size $n$ that is required for $\varepsilon$-accurate grid-based discretization with the mesh-size $h$,
(1) $|{\bf r}|\ll m\ll n=O(h^{-1}).$
The resulting Tucker rank is the most important parameter governing the complexity of subsequent multilinear algebra operations.
We present numerical experiments demonstrating the efficiency of the proposed techniques on nontrivial examples, including the computation of multi-particle electrostatic potentials arising in bio-molecular modeling and lattice type structures calculations in material science. In particular, it is gainfully applied to the long-range part of the many-particle electrostatic potentials in $\mathbb{R}^{3}$, obtained by the RS tensor format [6]. Numerical results indicate that the low-rank tensor representations can be computed with controllable accuracy, at a cost that depends only weakly on the number of particles [6]. In summary, we state the main benefits of the proposed method as follows:
  • Fast and accurate mesh-free decomposition: The method yields a quasi-optimal mesh-free Tucker decomposition for both regular functions and the long-range part of functions with singularities, such as the multi-particle electrostatic potential, by using rather small number of Chebyshev polynomials.
  • Efficiency for CP-formatted input: In case of input functions presented in the canonical (CP) tensor format, our method needs only a small number of 1D Chebyshev interpolations of skeleton vectors, followed by a reduced HOSVD [39, 35] of the Chebyshev coefficient tensor, thus resulting in a complexity that scales linear in the number of dimensions $d$.
  • Compatibility with grid-based methods: By discretizing the Chebyshev polynomials in the final ChebTuck format, one obtains a standard low-rank Tucker tensor on any desired fine grid. This avoids the direct, and often prohibitively expensive, decomposition of a large grid-based tensor in $\mathbb{R}^{n\times n\times n}$, while reproducing the results of established grid-based methods [39, 33].
The remainder of the paper is organized as follows. In Section 2, we formally define the hybrid Chebychev-Tucker tensor representation and discuss its transformation to the standard Tucker format. Section 3 presents the numerical schemes for constructing the approximation of trivariate functions, either given explicitly or only available on a regular grid, and provides corresponding error and complexity bounds. Section 4 first outlines the well-developed techniques for grid-based low-rank tensor representations of multi-particle interaction potentials. We then discuss in detail the Chebyshev-Tucker approximation of trivariate functions with singularities (the Newton kernel). We show efficiency of our methods for approximating the long-range part of the multi-particle electrostatic potentials of bio-molecular systems and lattice-type compounds of different sizes, and present various numerical tests illustrating the asymptotic performance of the proposed approach. Finally, Section 5 concludes the paper and outlines potential future research directions.
2 The mesh-free hybrid Chebychev-Tucker tensor format
In this section, we review the basics of tensors and introduce the hybrid Chebyshev-Tucker tensor format for the approximation of multivariate functions.
2.1 Definition of hybrid tensor format
Recall that a tensor of order $d$ is defined as a real multi-dimensional array over a $d$-tuple index set ${\bf A}=[a_{i_{1},\cdots,i_{d}}]\equiv[a(i_{1},\cdots,i_{d})]\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}},$ with multi-index notation ${\bf i}=(i_{1},\cdots,i_{d})$, $i_{\ell}=1,\cdots,n_{\ell}$. To get rid of the exponential scaling in storage size, one can apply the rank-structured formatted separable approximations of tensors. For example, the $R$-term canonical (CP) tensor format is defined by a finite sum of rank-1 tensors
(2) ${\bf U}_{R}={\sum}_{k=1}^{R}\xi_{k}{\bf u}_{k}^{(1)}\otimes\cdots\otimes{\bf u}_{k}^{(d)},\quad\xi_{k}\in\mathbb{R}_{+},$
where $R\in\mathbb{N}_{+}$ is the canonical rank, ${\bf u}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are normalized vectors, also known as the canonical (skeleton) vectors. The rank-$\bf r$}, ${\bf r}=(r_{1},\cdots,r_{d})$, orthogonal Tucker format is specified as the set of tensors ${\bf V}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ parametrized as
(3) ${\bf V}=\sum_{\nu_{1}=1}^{r_{1}}\cdots\sum_{\nu_{d}=1}^{r_{d}}\beta_{\nu_{1},\cdots,\nu_{d}}{\bf v}_{\nu_{1}}^{(1)}\otimes\cdots\otimes{\bf v}_{\nu_{d}}^{(d)}\equiv\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\cdots\times_{d}V^{(d)},$
where $\{{\bf v}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}}$ is a set of orthonormal vectors for $\ell=1,\cdots,d$. Here, $\times_{\ell}$ denotes the contraction along the $\ell$-th mode of the Tucker core $\bm{\beta}\in\mathbb{R}^{r_{1}\times\cdots\times r_{d}}$ with the second mode of the orthogonal matrices $V^{(\ell)}=[{\bf v}_{1}^{(\ell)},\cdots,{\bf v}_{r_{\ell}}^{(\ell)}]\in\mathbb{R}^{n_{\ell}\times r_{\ell}}$. Here and after we consider $d=3$ to simplify the presentation since (1) the extension to higher dimensions is straightforward, and (2) Tucker is not the recommended format for high dimensions due to the exponential scaling of the storage of the core tensor.
Next, we define the functional analog of the Tucker format and distinguish the discrete and functional two-level hybrid tensor formats in Definitions 2.1 and 2.2.
Definition 2.1 (Functional Tucker format).
We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $1,1]^{3}$ is represented in the functional rank-$\bf r$} Tucker format if it allows the factorization
(4) $f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}F^{(1)}(x_{1})\times_{2}F^{(2)}(x_{2})\times_{3}F^{(3)}(x_{3}),\quad F^{(\ell)}:[-1,1]\mapsto\mathbb{R}^{r_{\ell}},$
where $\bm{\beta}=[\beta_{\nu_{1},\nu_{2},\nu_{3}}]\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is the core tensor and $F^{(\ell)}(x_{\ell})$ is the Tucker factor function defined as a vector-valued function with functional entries, i.e., $F^{(\ell)}(x_{\ell})=(f_{1}^{(\ell)}(x_{\ell}),\cdots,f_{r_{\ell}}^{(\ell)}(x_{\ell}))$ for $f_{\nu_{\ell}}^{(\ell)}:[-1,1]\mapsto\mathbb{R}$, $\nu_{\ell}=1,\cdots,r_{\ell}$, $\ell=1,2,3$.
Definition 2.2 (Two-level hybrid Tucker formats).
(A) (Functional case) We say that $f$ is in the two-level hybrid functional rank-$({{\bf r}_{0},\bf r})$ Tucker format if the core tensor $\bm{\beta}$ in Eq. 4 is in turn a rank-${\bf r}_{0}$ Tucker tensor,
(5) $\bm{\beta}=\bm{\eta}\times_{1}G^{(1)}\times_{2}G^{(2)}\times_{3}G^{(3)},\quad G^{(\ell)}\in\mathbb{R}^{r_{\ell}\times r_{\ell}^{0}}.$
(B) (Algebraic case). We say that tensor ${\bf A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is parametrized in the two-level hybrid rank-$({{\bf r}_{0},\bf r})$ Tucker-Tucker algebraic format if it admits
(6) ${\bf A}=\bm{\beta}\times_{1}A^{(1)}\times_{2}A^{(2)}\times_{3}A^{(3)},\quad\mbox{with}\quad A^{(\ell)}=\{{\bf a}_{\nu_{\ell}}^{(\ell)}\}_{\nu_{\ell}=1}^{r_{\ell}}\in\mathbb{R}^{n_{\ell}\times r_{\ell}},\quad\ell=1,2,3,$
where the core coefficient $\bm{\beta}=\{\beta_{\nu_{1},\nu_{2},\nu_{3}}\}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is given by the rank-${\bf r}_{0}$ Tucker tensor with the small Tucker ranks $|{\bf r_{0}}|<|{\bf r}|$.
Furthermore, let $T_{k}(x):=\cos((k-1)\arccos(x))$ denote the (k1)(k-1)-st Chebyshev polynomial of the first kind, which forms a sequence of orthogonal polynomials on [1,1][-1,1] with respect to the weight $w(x)=1/\sqrt{1-x^{2}}$. If the functional entries of $F^{(\ell)}$ in Eq. 4 are specified by the Chebyshev polynomials, we say that $f$ is represented in the (two-level hybrid) Chebyshev-Tucker format, shortly ChebTuck format.
Definition 2.3 (ChebTuck format).
Let $T_{1:k}(x)=(T_{1}(x),\cdots,T_{k}(x))$ denote the vector-valued function of the first $k$ Chebyshev polynomials. We say that the real-valued function $f(x_{1},x_{2},x_{3})$ defined on $1,1]^{3}$ is represented in the ChebTuck-$(\bf{m},\bf{r})$ format if it allows the factorization
(7) $f(x_{1},x_{2},x_{3})={\bf C}\times_{1}T_{1:m_{1}}(x_{1})\times_{2}T_{1:m_{2}}(x_{2})\times_{3}T_{1:m_{3}}(x_{3}),$
where ${\bf C}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ is a rank-$\bf{r}$ Tucker tensor given by
(8) ${\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}},\quad V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}.$
For the practical application in scientific computing the presented hybrid Tucker-type tensor representation might be easily transformed to the standard Tucker-type format when required.
2.2 Transforming hybrid representation to the standard Tucker format
The following Proposition 2.4 proves that any tensor parametrized in the two-level orthogonal Tucker format with ranks $({\bf r}_{0},{\bf r})$ can be converted to the standard Tucker tensor with an reduced (quasi-optimal) rank $|{\bf r}_{0}|<|{\bf r}|$. This allows the traditional multilinear algebra operations with tensors in the hybrid format.
Proposition 2.4.
Let the two-level orthogonal Tucker tensor ${\bf A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be given in the form Eq. 6, where the core $\bm{\beta}$ is parametrized by the rank ${\bf r}^{0}=(r_{1}^{0},r_{2}^{0},r_{3}^{0})$ orthogonal Tucker tensor as in Eq. 5. Then $\bf A$} takes the form of a rank ${\bf r}^{0}$ orthogonal Tucker representation with the core tensor $\bm{\eta}\in\mathbb{R}^{r_{1}^{0}\times r_{2}^{0}\times r_{3}^{0}}$ and the orthogonal Tucker factor matrices $A^{(\ell)}G^{(\ell)}\in\mathbb{R}^{n_{\ell}\times r_{\ell}^{0}}$, i.e.,
(9) ${\bf A}=\bm{\eta}\times_{1}(A^{(1)}G^{(1)})\times_{2}(A^{(2)}G^{(2)})\times_{3}(A^{(3)}G^{(3)}).$
Proof 2.5.
Substituting $\bm{\beta}$ Eq. 5 into Eq. 6 and applying contracted product operations to the resultant Tucker tensor proves Eq. 9. The orthogonality of the factor matrices follows from $(G^{(\ell)})^{\top}(A^{(\ell)})^{\top}A^{(\ell)}G^{(\ell)}=I_{r_{\ell}^{0}}$.
Notice that the resultant conventional Tucker tensor (9) can be easily transformed to the canonical tensor format by using the so-called Tuck-to-can transform applied to the core tensor $\bm{\eta}$, see [39, Def. 2.3, (2.6)], [38, Rem. 2.7], and [34].
The following statement is important in further use of the hybrid functional representation in multilinear algebra, see also (17).
Proposition 2.6.
The format Eq. 7 can be converted to the standard functional Tucker format as in Definition 2.1, i.e., Eq. 7 can be equivalently written as
(10) $f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}),$
where the Tucker factor functions are given by $v^{(\ell)}(x_{\ell})=T_{1:m_{\ell}}(x_{\ell})V^{(\ell)}\in\mathbb{R}^{r_{\ell}}$, i.e., each entry of $v^{(\ell)}$ is a linear combination of the Chebyshev polynomials, while the core tensor is given by Eq. 8. Furthermore, if Eq. 8 is an orthogonal Tucker tensor, then $\{T_{1:m_{\ell}}(x_{\ell})V^{(\ell)}\}$ is a set of $r_{\ell}^{0}$ orthogonal polynomials with the inner product induced by the orthogonality property of Chebyshev polynomials.
The proof is in essence the same as Proposition 2.4, i.e., follows from the easy calculation of contracted products. We nevertheless write out several useful equivalent representations of the ChebTuck format explicitly for further use.
Proof 2.7.
From Eq. 7 and Eq. 8, we have
(11) $f(x_{1},x_{2},x_{3})=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})$
$=$ $\sum_{j_{1}=1}^{r_{1}}\sum_{j_{2}=1}^{r_{2}}\sum_{j_{3}=1}^{r_{3}}\beta_{j_{1},j_{2},j_{3}}\sum_{i_{1}=1}^{m_{1}}V^{(1)}_{i_{1},j_{1}}T_{i_{1}}(x_{1})\sum_{i_{2}=1}^{m_{2}}V^{(2)}_{i_{2},j_{2}}T_{i_{2}}(x_{2})\sum_{i_{3}=1}^{m_{3}}V^{(3)}_{i_{3},j_{3}}T_{i_{3}}(x_{3})$
$=$ $\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}),$
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $\ell=1,2,3$. The desired result then follows. The orthogonality property is based on the argument analogues to that of the algebraic case but reformulated for polynomial arrays. We note that the inner product should be properly reweighted.
Remark 2.8.
We can see that the ChebTuck format is a degree ${\bf m}=(m_{1},m_{2},m_{3})$ polynomial in Chebyshev basis, where the coefficients are given as a Tucker tensor. For this reason, we sometimes also write a subscript $\bf m$} to denote the ChebTuck format of certain degrees, e.g., $f_{\bf m}$.
3 Numerical methods for hybrid Chebyshev-Tucker approximation of trivariate functions
In what follows, we introduce the numerical methods for rank-structured approximation of trivariate functions in the novel ChebTuck tensor format introduced in Definition 2.3. The basic idea is to first compute the Chebyshev coefficient tensor $\bf C$}, either in full tensor format or in some low-rank format, and then convert $\bf C$} to the desired Tucker format with quasi-optimal $\varepsilon$-rank. Our techniques can be applied directly to multivariate functions with moderate regularity for which the Chebyshev tensor product interpolation demonstrates satisfactory convergence rates. Furthermore, we extend the presented approach beyond the class of regular functions and successfully apply it to functions with multiple singularities by using the range-separated tensor decomposition.
3.1 Classical multivariate Chebyshev interpolant
3.1.1 Functional case
Given $f:[-1,1]^{3}\to\mathbb{R}$, we consider its approximation by the tensor product of Chebyshev polynomials in the form (c.f. Eq. 7)
(12) $f(x_{1},x_{2},x_{3})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3}),$
where $T_{k}(x)$ is the (k1)(k-1)-st Chebyshev polynomial, ${\bf C}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ is the Chebyshev coefficient tensor (CCT).
There are mainly two ways to view the approximant $\tilde{f}_{\bf m}$ which lead to two different methods to compute the CCT $\bf C$}. The first way is to view $\tilde{f}_{\bf m}$ as a truncation of the tensor product Chebyshev series expansion of $f$ and compute the coefficients by projection and integration. In this case, the Chebyshev coefficients are in fact straightforward to derive and given by the $w$-weighted inner product of $f$ and the corresponding Chebyshev basis, i.e., ${\bf C}_{i_{1},i_{2},i_{3}}=\int_{[-1,1]^{3}}f(\mathbf{x})T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})w(\mathbf{x})\operatorname{d}\mathbf{x}$, where $w(\mathbf{x})=\prod_{\ell=1}^{3}(1-x_{\ell}^{2})^{-1/2}$ is the Chebyshev weight function, see e.g., [45, 46], but the computation is troublesome since one needs to evaluate those 3D integrals. Here, we used the notation $\mathbf{x}=(x_{1},x_{2},x_{3})$. The second way is to view $\tilde{f}_{\bf m}$ as an interpolation of $f$ at Chebyshev nodes, namely, $\tilde{f}_{\bf m}$ satisfies the interpolatory property $\tilde{f}_{\bf m}(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})$ for all $i_{\ell}=1,\ldots,m_{\ell}$, $\ell=1,2,3$, where $s_{i_{\ell}}^{(\ell)}$ are the Chebyshev nodes of the second kind on each mode given by $s_{i_{\ell}}^{(\ell)}=\cos\left((i_{\ell}-1)\pi/(m_{\ell}-1)\right)$. The derivation of the CCT $\bf C$} in this case is more involved while it turns out that the computation can be done efficiently by the discrete cosine transform (DCT). The accuracy of these two approaches is in practice similar [45] and we will therefore focus on the interpolation approach in this paper.
For this purpose, we first define the function evaluation tensor ${\bf T}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ of $f$ at Chebyshev nodes as
(13) ${\bf T}_{i_{1},i_{2},i_{3}}=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)}),\quad i_{\ell}=1,\ldots,m_{\ell},\quad\ell=1,2,3$
and the inverse DCT matrix $W^{(\ell)}\in\mathbb{R}^{m_{\ell}\times m_{\ell}}$ as
(14) $W^{(\ell)}=\dfrac{2}{m_{\ell}-1}\begin{pmatrix}\frac{1}{4}T_{1}(s_{1})&\frac{1}{2}T_{1}(s_{2})&\frac{1}{2}T_{1}(s_{3})&\dots&\frac{1}{4}T_{1}(s_{m_{\ell}})\\ \frac{1}{2}T_{2}(s_{1})&T_{2}(s_{2})&T_{2}(s_{3})&\dots&\frac{1}{2}T_{2}(s_{m_{\ell}})\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{4}T_{m_{\ell}}(s_{1})&\frac{1}{2}T_{m_{\ell}}(s_{2})&\frac{1}{2}T_{m_{\ell}}(s_{3})&\dots&\frac{1}{4}T_{m_{\ell}}(s_{m_{\ell}})\end{pmatrix}$,
then by [47, Equations 6.27—6.28], the CCT $\bf C$} can be computed by the DCT of the function evaluation tensor $\bf T$} as
(15) ${\bf C}={\bf T}\times_{1}W^{(1)}\times_{2}W^{(2)}\times_{3}W^{(3)}.$
Given an error tolerance $\varepsilon>0$, we can approximate the coefficient tensor $\bf C$} by a Tucker tensor with almost optimal rank parameter ${\bf r}=(r_{1},r_{2},r_{3})$ using the Tucker-ALS algorithm denoted by $\texttt{tuck\_als}({\bf C},\varepsilon)$
(16) ${\bf C}\approx\hat{\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad[\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$
with $\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ and $V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}$, $\ell=1,2,3$. See Remark 3.1 for the discussions on Tucker-ALS and other Tucker decomposition algorithms. Substituting $\bf C$} in Eq. 12 by its Tucker format Eq. 16 gives the ChebTuck approximation of $f$, denoted by $\hat{f}_{\bf m}$. In view of Proposition 2.6, the ChebTuck approximation $\hat{f}$ of $f$ can be written as
(17) $f(\textbf{x})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})\approx\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}).$
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $j_{\ell}=1,\ldots,r_{\ell}$, $\ell=1,2,3$.
In summary, this leads to the following straightforward Algorithm 1 for constructing the Chebyshev-Tucker approximation of a function $f:[-1,1]^{3}\to\mathbb{R}$.
Algorithm 1 ChebTuck Approximation: functional case
1:Given a function $f:[-1,1]^{3}\to\mathbb{R}$, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and tuck_als tolerance $\varepsilon>0$.
2:The ChebTuck format $\hat{f}_{\bf m}$ of $f$.
3:Compute the function evaluation tensor $\bf T$} by Eq. 13, then the CCT $\bf C$} by Eq. 15.
4:Compute the Tucker decomposition: $\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$ .
5:Form the factor functions $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $\ell=1,2,3$.
6:Return $\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3})$.
Remark 3.1.
Notice that tuck_als in line 4 of Algorithm 1 implementing the classical Tucker-ALS algorithm [18, 19] for this purpose, is widely used in practice for moderate $d$ and $m$. It enjoys strong theoretical guarantees such that for any given $\varepsilon>0$, it approximates the original tensor in the sense of relative error $\|{\bf C}-\hat{\bf C}\|_{\mathrm{F}}\leq\varepsilon\|{\bf C}\|_{\mathrm{F}}$ with almost optimal rank parameters, where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm of a tensor. However, it is computational expensive since it requires the access to all entries of the tensor $\bf C$}, amounting to $O(m^{3})$ space complexity and successive SVDs of the unfolding matrices, amounting to $O(m^{4})$ algebraic operations. Alternatives to the classical Tucker-ALS are, e.g., the multigrid algorithm [39], cross-based algorithms [49, 21, 20, 50, 29], or randomized algorithms [15, 40]. In case of the relatively small rank of the input tensor, those algorithms are more efficient in terms of computational cost and memory usage. In particular, using the cross-based algorithms can avoid the evaluation of the full tensor $\bf T$}. However, they contain several heuristics, especially the cross-based algorithms have no theoretical accuracy guarantees for given $\varepsilon$. The discussions on the difference and choice of Tucker decomposition algorithms are beyond the scope of this paper. We use the classical Tucker-ALS algorithm in our theoretical analysis and numerical experiments demonstrated reliable efficiency for low-rank approximation of the coefficient tensor $\bf C$}.
3.1.2 Algebraic case with full tensor input
In many applications, see [37, 34, 7] among others, including ours on multi-particle modeling which will be discussed in Section 4, the function $f$ may not be given explicitly in the full domain $1,1]^{3}$, but only the data on the equi-spaced grid are available. Specifically, we only have access to the function values at uniform grid points collected in a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, ${\bf F}_{i_{1},i_{2},i_{3}}=f(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})$ for $i_{\ell}=1,\ldots,n_{\ell}$, $\ell=1,2,3$. Here $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ is the grid point and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size.
In this case, we first extrapolate the function to the full domain $1,1]^{3}$ by constructing a multivariate cubic spline interpolation $Q(x_{1},x_{2},x_{3})$ of the data $\bf F$} on the uniform grid, i.e., $Q(x_{1},x_{2},x_{3})$ is a $C^{2}$ piecewise cubic polynomial satisfying the interpolatory property $Q(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})={\bf F}_{i_{1},i_{2},i_{3}}$, and then call Algorithm 1 with $f=Q$. This procedure is summarized in Algorithm 2. Note that here and forehand, we assume $n_{\ell}\geq 4$ naturally to ensure the existence of the cubic spline.
Algorithm 2 ChebTuck Approximation: algebraic case
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathbf{m}$, $\varepsilon>0$ as in Algorithm 1.
2:The ChebTuck format $\hat{f}_{\bf m}$ of $f$.
3:Construct the multivariate cubic spline interpolation $Q(x_{1},x_{2},x_{3})$ of the data $\bf F$} on the uniform grid by, e.g., interpn in MATLAB.
4:Call Algorithm 1 with $f=Q$.
3.1.3 Algebraic case with CP tensor input
In fact, the multivariate cubic spline interpolation, see e.g. [17, Chapter 17], is computationally expensive. However, in many applications of interest, e.g., quantum chemistry [34], multi-particle modeling [37] and the discretization of certain differential operators [32], the target input tensor $\bf F$} has some inherent low-rank structure and thus can be cheaply represented in some low-rank tensor format. In such cases, the computation of the evaluation tensor $\bf T$} can be done more efficiently. In particular, we consider here the case where $\bf F$} is given in the CP format, i.e.,
(18) ${\bf F}={\sum}_{k=1}^{R}\xi_{k}{\bf a}_{k}^{(1)}\otimes{\bf a}_{k}^{(2)}\otimes{\bf a}_{k}^{(3)},$
where ${\bf a}_{k}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are $\ell_{2}$ normalized canonical vectors of each mode. In this case, we only need to construct univariate cubic splines of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode, and make tensor products and summations of them. Specifically, let $q_{k}^{(\ell)}(x_{\ell})$ be the cubic spline of the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$, i.e., $q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})={\bf a}_{k}^{(\ell)}(i_{\ell})$ and then construct the sum of the tensor product of the univariate splines as
(19) $Q(x_{1},x_{2},x_{3})={\sum}_{k=1}^{R}\xi_{k}q_{k}^{(1)}(x_{1})q_{k}^{(2)}(x_{2})q_{k}^{(3)}(x_{3}).$
Calling Algorithm 1 with $f=Q$, one sees immediately from the definitions of the CCT $\bf C$} Eqs. 13 and 15 that $\bf C$} computed in line 3 is also in CP format. Specifically, let ${\bf q}_{k}^{(\ell)}\in\mathbb{R}^{m_{\ell}}$ contain the function evaluations of the univariate spline $q_{k}^{(\ell)}(x)$ at $\{s_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{m_{\ell}}$, i.e., ${\bf q}_{k}^{(\ell)}(i_{\ell})=q_{k}^{(\ell)}(s_{i_{\ell}}^{(\ell)})$, then by definition of the function evaluation tensor $\bf T$} in Eq. 13 and the CCT $\bf C$} in Eq. 15, we have
(20) ${\bf T}={\sum}_{k=1}^{R}\xi_{k}{\bf q}_{k}^{(1)}\otimes{\bf q}_{k}^{(2)}\otimes{\bf q}_{k}^{(3)}\implies{\bf C}={\sum}_{k=1}^{R}\xi_{k}{\bf c}_{k}^{(1)}\otimes{\bf c}_{k}^{(2)}\otimes{\bf c}_{k}^{(3)},$
with ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ for $\ell=1,2,3$ and $W^{(\ell)}$ being the inverse DCT matrix defined in Eq. 14. From here, we can leverage the reduced HOSVD (rhosvd) [39] method instead of full tuck_als to compress $\bf C$} directly from canonical to Tucker format, without forming the full tensor $\bf C$}. The rest of the algorithm remains the same as Algorithm 1. This procedure is summarized in Algorithm 3.
Algorithm 3 ChebTuck Approximation: algebraic case (CP tensor input)
1:Given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ in CP format, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and rhosvd tolerance $\varepsilon>0$.
2:The ChebTuck format $\hat{f}_{\bf m}$ of $f$.
3:Compute the splines $q_{k}^{(\ell)}(x_{\ell})$ and the evaluations ${\bf q}_{k}^{(\ell)}$ for $k=1,\ldots,R$, $\ell=1,2,3$.
4:Compute the canonical vectors ${\bf c}_{k}^{(\ell)}=W^{(\ell)}{\bf q}_{k}^{(\ell)}$ of $\bf C$} for $k=1,\ldots,R$, $\ell=1,2,3$.
5:Compute the reduced Tucker decomposition: $\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{rhosvd}({\bf C},\varepsilon)$
6:Perform lines 5-6 in Algorithm 1.
3.2 Error bounds for the Chebyshev-Tucker approximation
In this section, we always assume $m=m_{1}=m_{2}=m_{3}$ without loss of generality to simplify the notation. While we state the results for general tensor order $d$ since they also hold for $d\geq 3$, the proofs specialize to $d=3$ for clarity of presentation.
3.2.1 Functional case
We first consider the ChebTuck approximation $\hat{f}_{\bf m}$ of a function $f$ obtained by Algorithm 1, i.e. the functional case. It is well-known that multivariate Chebyshev interpolation $\tilde{f}_{\bf m}$ converges exponentially fast to the function $f$ in the uniform norm over $1,1]^{3}$ with respect to the degree $\bf m$} of the Chebyshev polynomials, provided that $f$ has sufficient regularity. Our ChebTuck format introduces an additional error due to the Tucker approximation of the Chebyshev coefficient tensor $\bf C$}, which is controlled by the Tucker-ALS truncation error $\varepsilon$. In the following proposition, we show that the ChebTuck approximation $\hat{f}_{\bf m}$ is almost as accurate as the full Chebyshev interpolation $\tilde{f}_{\bf m}$ provided that the Tucker approximation error $\varepsilon$ is sufficiently small.
Proposition 3.2.
Given notations developed above, we have
(21) $\|f-\hat{f}_{\bf m}\|_{\infty}\leq\|f-\tilde{f}_{\bf m}\|_{\infty}+m^{d/2}\varepsilon\|\bf{C}\|_{\mathrm{F}},$
where $\|\cdot\|_{\infty}$ is the uniform norm over $1,1]^{3}$.
Proof 3.3.
By triangle inequality, it suffices to prove $\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq m^{3/2}\varepsilon\|\bf{C}\|_{\mathrm{F}}$. Indeed, by the definition of the ChebTuck format Eq. 17 and the fact that Chebyshev polynomials are bounded by 1, we have
$\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}$ $=\left\|\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{m}\sum_{i_{3}=1}^{m}({\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}})T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})\right\|_{\infty}$
$\leq\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{m}\sum_{i_{3}=1}^{m}|{\bf C}_{i_{1},i_{2},i_{3}}-\hat{\bf C}_{i_{1},i_{2},i_{3}}|\leq m^{3/2}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}\leq m^{3/2}\varepsilon\|{\bf C}\|_{\mathrm{F}},$
where we used the Cauchy-Schwarz inequality and the relative truncation error $\varepsilon$ of the Tucker-ALS approximation in the last two inequalities.
3.2.2 Algebraic case
Now, we consider the ChebTuck approximation in the algebraic case, i.e., given a tensor ${\bf F}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ assumed to discretize a black-box function $f$ on the uniform grid. In this case, the metric for the error of the approximation is the difference between the ChebTuck approximation $\hat{f}_{\bf m}$ evaluated at the uniform grid and the tensor $\bf F$}, i.e., $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|$. In other words, we hope to recover the original input $\bf F$} when we evaluate the ChebTuck format $\hat{f}_{\bf m}$ at the uniform grid. Since we know nothing about the function $f$ except for the values on the grid, for Algorithm 2, we can say nothing about the error of the Chebyshev interpolation of the multivariate cubic spline $Q$, thus neither the ChebTuck approximation $\hat{f}_{\bf m}$. For example, for highly non-regular function $f$ we would not expect the Chebyshev interpolation thus also the ChebTuck approximation to be accurate.
Fortunately, in the case of CP tensor input, we can provide a bound for the error of the ChebTuck approximation $\hat{f}_{\bf m}$ depending on the errors of the univariate Chebyshev interpolations of the canonical vectors ${\bf a}_{k}^{(\ell)}$ in each mode. Note that the canonical vectors ${\bf c}_{k}^{(\ell)}$ of the Chebyshev coefficient tensor $\bf C$} are nothing but the univariate Chebyshev coefficients of the splines $q_{k}^{(\ell)}(x)$. In other words, let $g_{k}^{(\ell)}(x):=\sum_{j=1}^{m}{\bf c}_{k}^{(\ell)}(j)T_{j}(x)$, then $g_{k}^{(\ell)}(x)$ is the Chebyshev interpolation of $q_{k}^{(\ell)}(x)$. Let $\hat{\bf C}$ be the RHOSVD truncation of $\bf C$}, then the ChebTuck approximation error is given by
(22) $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|+\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}$
where $\tilde{f}_{\bf m}$ is the Chebyshev interpolant Eq. 12 with coefficients $\bf C$} given in Eq. 20, i.e., before RHOSVD truncation. $\tilde{f}_{\bf m}$ is thus given as
(23) $\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})={\sum}_{k=1}^{R}\xi_{k}g_{k}^{(1)}(x_{1})g_{k}^{(2)}(x_{2})g_{k}^{(3)}(x_{3}).$
Then we have the following bound for the first term in the right-hand side of Eq. 22.
Lemma 3.4.
Let $\tilde{f}_{\bf m}$ be of the form Eq. 12 with coefficients $\bf C$} in Eq. 20, i.e., $\tilde{f}_{\bf m}$ takes the form Eq. 23. Let $\delta_{k}^{(\ell)}:=\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}$ be the error of the univariate Chebyshev interpolation of $q_{k}^{(\ell)}(x)$ and $\delta:=\max_{k,\ell}\delta_{k}^{(\ell)}$, then we have
(24) $\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(d)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\left((1+\delta)^{d}-1\right),$
where $\|\bm{\xi}\|_{1}=\sum_{k=1}^{R}|\xi_{k}|$ is the $\ell_{1}$ norm of the normalization constants.
Proof 3.5.
Recall that by construction, $q_{k}^{(\ell)}$ interpolates the data ${\bf a}_{k}^{(\ell)}$ on the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$ and the Chebyshev interpolants have uniform error $\delta_{k}^{(\ell)}$. Therefore, for each $\ell=1,\ldots,d$, $k=1,\ldots,R$ and $i_{\ell}=1,\ldots,n_{\ell}$, we have
(25) $|{\bf a}_{k}^{(\ell)}(i_{\ell})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|=|q_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq\|q_{k}^{(\ell)}-g_{k}^{(\ell)}\|_{\infty}=\delta_{k}^{(\ell)}$
and in particular,
(26) $|g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})|\leq|{\bf a}_{k}^{(\ell)}(i_{\ell})|+\delta_{k}^{(\ell)}\leq\|{\bf a}_{k}^{(\ell)}\|_{\infty}+\delta_{k}^{(\ell)}\leq 1+\delta_{k}^{(\ell)},$
where we have used the fact that the canonical vectors are $\ell_{2}$ normalized. Therefore, for any $i_{1},\ldots,i_{d}$, by the definition of $\tilde{f}_{\bf m}$ in Eq. 23 and $\bf F$} in Eq. 18, we have
$\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},\ldots,t_{i_{d}}^{(d)})-{\bf F}_{i_{1},\ldots,i_{d}}\right|\leq{\sum}_{k=1}^{R}\xi_{k}\left|\prod_{\ell=1}^{d}g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-\prod_{\ell=1}^{d}{\bf a}_{k}^{(\ell)}(i_{\ell})\right|$
$\leq$ $\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\left|\prod_{r=1}^{\ell-1}g_{k}^{(r)}(t_{i_{\ell}}^{(r)})\right|\left|g_{k}^{(\ell)}(t_{i_{\ell}}^{(\ell)})-{\bf a}_{k}^{(\ell)}(i_{\ell})\right|\left|\prod_{r=\ell+1}^{d}{\bf a}_{k}^{(r)}(i_{r})\right|$
$\leq$ $\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\prod_{r=1}^{\ell-1}\left(1+\delta_{k}^{(r)}\right)\delta_{k}^{(\ell)}\leq\sum_{k=1}^{R}\xi_{k}\sum_{\ell=1}^{d}\left(1+\delta\right)^{\ell-1}\delta=\|\bm{\xi}\|_{1}\left((1+\delta)^{d}-1\right),$
where the second inequality results from applying a standard telescopic argument.
By assuming $\delta<1/d$, which is easily satisfied in practical computations, we can further improve the bound Eq. 24 to a linear growth with respect to the dimension $d$ as in the following corollary.
Corollary 3.6.
Under the assumptions of Lemma 3.4, further assuming $\delta d<1$ yields
(27) $\max_{i_{1},i_{2},i_{3}}\left|\tilde{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\min\left\{\dfrac{\delta d}{1-\delta d},(\mathrm{e}-1)\delta d\right\}.$
Proof 3.7.
First note that $(1+\delta)^{d}\leq\mathrm{e}^{\delta d}$. Then the desired result follows from the elementary inequalities $\mathrm{e}^{x}\leq(\mathrm{e}-1)x+1\text{ and }\mathrm{e}^{x}\leq(1-x)^{-1}$ for $x\in(0,1)$.
The following lemma states that the error of the univariate Chebyshev interpolation of the cubic spline $q_{k}^{(\ell)}$ decays cubically with respect to the Chebyshev degree $m$.
Lemma 3.8.
Let $V_{k}^{\ell}<+\infty$ be the bounded variation of the third derivative of the spline $q_{k}^{(\ell)}$ and $V=\max_{k,\ell}V_{k}^{\ell}$. Then we have
(28) $\delta_{k}^{(\ell)}\leq\dfrac{4V_{k}^{\ell}}{3\pi(m-3)^{3}}\text{ for }k=1,\ldots,R,\quad\ell=1,2,3\text{ and }\delta\leq\dfrac{4V}{3\pi(m-3)^{3}}.$
Proof 3.9.
By definition, $q_{k}^{(\ell)}$ is a $C^{2}$ piecewise cubic polynomial, and thus absolutely continuous differentiable until the second derivative. The third derivative is piecewise constant thus of bounded variation $V_{k}^{\ell}$. Note that $V_{k}^{\ell}$ is uniquely determined by the spline knots (here the uniform grid $\{t_{i_{\ell}}^{(\ell)}\}_{i_{\ell}=1}^{n_{\ell}}$) and the spline values ${\bf a}_{k}^{(\ell)}(i_{\ell})$. The results then follow from the standard error bound for Chebyshev interpolation of differentiable functions, see e.g., [59, Thm. 7.2].
Now, we consider the error introduced by the RHOSVD truncation [39] in Algorithm 3, namely, the second term in the right-hand side of Eq. 22.
Lemma 3.10.
Let $\sigma_{k}^{(\ell)}$, $k=1,\ldots,\min\{m,R\}$ be the descending singular values of the side matrices $C^{(\ell)}=[{\bf c}_{1}^{(\ell)},\cdots,{\bf c}_{R}^{(\ell)}]\in\mathbb{R}^{m\times R}$ of the Tucker format Eq. 20, and $r_{\ell}$ be the truncation rank of the RHOSVD approximation, then we have
(29) $\|\hat{f}_{\bf m}-\tilde{f}_{\bf m}\|_{\infty}\leq m^{d/2}\|\bm{\xi}\|\sum_{\ell=1}^{d}\Big(\sum_{k=r_{\ell}+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2},\text{ where }\|\bm{\xi}\|^{2}=\sum_{k=1}^{R}\xi_{k}^{2}.$
Proof 3.11.
By similar arguments as in the proof of Proposition 3.2, we have $\|\tilde{f}_{\bf m}-\hat{f}_{\bf m}\|_{\infty}\leq m^{d/2}\|{\bf C}-\hat{\bf{C}}\|_{\mathrm{F}}$. The result then follows from the standard error analysis of the RHOSVD truncation [39, Thm. 2.5]. See also [35] for a more recent presentation.
Combining Corollary 3.6, Lemma 3.10 and the triangle inequality Eq. 22, we arrive at the following theorem immediately, which provides a bound for the error of the ChebTuck approximation in the algebraic case of CP tensor input.
Theorem 3.12.
With the notations above, the error of the ChebTuck approximation $\hat{f}_{\bf m}$ of a CP tensor $\bf F$} produced by Algorithm 3 is bounded by
$\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\leq\|\bm{\xi}\|_{1}\mathrm{e}\delta d+m^{d/2}\|\bm{\xi}\|\sum_{\ell=1}^{d}\Big(\sum_{k=r_{\ell}+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2}.$
Remark 3.13.
From the above theorem, one can design an adaptive algorithm that constructs the ChebTuck approximation of a CP tensor with a given error $\varepsilon>0$ by choosing the smallest possible $r_{\ell}$ in each mode $\ell=1,2,3$ such that the error bound is less than $\varepsilon$. This is because all the terms in the bound are computable. Specifically, $\delta$ will be anyway computed when we construct the univariate Chebyshev interpolants, and the singular values $\sigma_{k}^{(\ell)}$ are available after the SVD decomposition of the side matrices in the CP format.
3.2.3 The case of multi-particle potentials
In this section, we consider the ChebTuck approximation error for multi-particle potentials of the form Eq. 33, which is also the main application area in our numerical experiments. We show that with the use of the range-separated tensor format [6], the error of the ChebTuck approximation of the long-range part decays exponentially fast with respect to the truncation rank $r_{\ell}$, i.e., we have the following Corollary 3.14 for the second term of the error bound in Theorem 3.12.
Corollary 3.14.
Suppose ${\bf F}\in\mathbb{R}^{n\times n\times n}$ is the tensor representation of the long-range part of an $N$-particle potential satisfying the conditions of [6, Thm. 3.1]. Then the singular values of the side matrices $C^{(\ell)}$ decay exponentially fast, resulting in the following error bound
(30) $\sum_{\ell=1}^{d}\Big(\sum_{k=r+1}^{\min\{m,R\}}(\sigma_{k}^{(\ell)})^{2}\Big)^{1/2}\leq C_{m,n}Nd\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right),$
where, given the target function $f$, the constant $C=C_{m,n}$ depends only on the Chebyshev degree $m$ and the size $n$ of the uniform grid. In particular, the $\varepsilon$-Tucker rank of the canonical CCT tensor $\bf C$} is estimated by $\mathcal{O}(\log^{3/2}|\log(\varepsilon/C_{m,n}N)|)$.
Proof 3.15.
Let $A^{(\ell)}=[{\bf a}_{1}^{(\ell)},\cdots,{\bf a}_{R}^{(\ell)}]\in\mathbb{R}^{n\times R}$ be the side matrix of the CP tensor $\bf F$} Eq. 18. Then [6, Thm. 3.1] implies that the $\varepsilon$-rank of the side matrix $A^{(\ell)}$ has the bound $r_{\varepsilon}(A^{\ell})\leq C\log^{3/2}(|\log(\varepsilon/N)|)$ thus we have
(31) $\Big(\sum_{k=r_{\ell}+1}^{\min\{n,R\}}(\sigma_{k}^{(\ell)}(A^{(\ell)}))^{2}\Big)^{1/2}\leq CN\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right).$
Let $P^{(\ell)}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, ${\bf q}_{k}^{(\ell)}=P^{(\ell)}{\bf a}_{k}^{(\ell)}$ be the operator that maps the canonical vector ${\bf a}_{k}^{(\ell)}$ to the evaluations of the spline $q_{k}^{(\ell)}$ at the Chebyshev points, i.e., the procedure in line 3 of Algorithm 3. It can be easily verified that the $P^{(\ell)}$ is linear thus can be represented as a matrix in $\mathbb{R}^{m\times n}$. Therefore, we have ${\bf c}_{k}^{(\ell)}=W^{(\ell)}P^{(\ell)}{\bf a}_{k}^{(\ell)}$ by line 4 of Algorithm 3, hence ${C}^{(\ell)}=W^{(\ell)}P^{(\ell)}A^{(\ell)}$. Therefore we have $\sigma_{k}^{(\ell)}\leq\|W^{(\ell)}P^{(\ell)}\|\sigma_{k}^{(\ell)}(A^{(\ell)})$ for $k=1,\ldots,R$ and $\ell=1,2,3$ by the min-max characterization of singular values. It is clear that $\|W^{(\ell)}P^{(\ell)}\|$ only depends on $m$ and $n$ since the spline knots and the Chebyshev points are only defined by $m$ and $n$ in the fixed interval [1,1][-1,1]. Absorbing the constant $C$ and the factor $\|W^{(\ell)}P^{(\ell)}\|$ into $C_{m,n}$ completes the proof.
Remark 3.16.
Regarding the constant $C_{m,n}$, numerical estimates show that it is in fact decaying polynomially with respect to both $m$ and $n$, though we do not have a proof for this claim. Combining Corollary 3.14 with Theorem 3.12 and Lemma 3.8, we can conclude that for the long-range part of the $N$-particle potential, the ChebTuck approximation error has the following estimate
(32) $\max_{i_{1},i_{2},i_{3}}\left|\hat{f}_{\bf m}(t_{i_{1}}^{(1)},t_{i_{2}}^{(2)},t_{i_{3}}^{(3)})-{\bf F}_{i_{1},i_{2},i_{3}}\right|\lesssim dm^{-3}+m^{d/2}Nd\exp\left(-\exp\left(r^{\frac{2}{3}}\right)\right).$
For fixed dimension $d$, in particular for practical applications to multi-particle potentials where $d=3$, the doubly exponential decay rate in the truncation rank $r$ essentially compensates the growth $m^{d/2}$ in the second term, resulting in a very favorable error bound. The results are consistent with our numerical experiments in Section 4.2.
3.3 Complexity bounds for Algorithms 1, 2, and 3
In this section, we briefly discuss the time complexity of the proposed algorithms for the ChebTuck approximation of a function in both functional and algebraic cases.
Proposition 3.17.
For the functional case Algorithm 1, let $C_{f}$ be the cost of one functional call. For the algebraic cases Algorithms 2 and 3, let $n=\max_{\ell}n_{\ell}$ be the maximum grid size. For the algebraic cases with CP tensor input Algorithms 3 and 3, let $R$ be the rank of the CP tensor. Let $m$ be Chebyshev degree in each mode and $\varepsilon>0$ be the Tucker-ALS or RHOSVD tolerance. Then the complexity of the algorithms is estimated by
Algorithm 1: $C_{f}m^{3}+C_{1}m^{3}\log m+C_{2}m^{4}+C_{3}m^{4}|\log\varepsilon|$.
Algorithm 2: $C_{4}m^{3}+C_{5}n^{3}+C_{1}m^{3}\log m+C_{2}m^{4}+C_{3}m^{4}|\log\varepsilon|$.
Algorithm 3: $C_{5}nR+C_{6}mR+C_{7}mR\log m+C_{8}mR\min\{m,R\}+C_{9}R|\log\varepsilon|^{3}$.
Proof 3.18.
For the functional case Algorithm 1, the number of functional calls is $m^{3}$ and therefore the cost of evaluating $\bf T$} is $C_{f}m^{3}$. The cost of computing the CCT $\bf C$} is $C_{1}m^{3}\log m$ by the FFT-based Chebyshev coefficient computation. The cost of Tucker-ALS is $C_{2}m^{4}+C_{3}m^{4}r$ by the standard results on its complexity, see e.g. [37, Chapter 3.4.3], where the Tucker rank $r=\mathcal{O}(|\log\varepsilon|)$ is assumed. This assumption is true for a wide range of functions including Newton kernel in our numerical experiments [38]. This completes the proof for the functional case.
For the algebraic cases, we first claim that the cost of evaluating a cubic spline interpolation at $m$ points given data on uniform $n$-point grid is $\mathcal{O}(m+n)$. Indeed, the computation of the spline coefficients requires solving a tridiagonal system of size 4n44n-4, which costs $\mathcal{O}(n)$, and an evaluation of the spline at $m$ points costs $\mathcal{O}(m)$ in case the $n$-point grid is uniform. Similarly, the cost of trivariate cubic spline interpolation which is required in Algorithm 2 is $\mathcal{O}(m^{3}+n^{3})$. This then completes the proof for the complexity of Algorithm 2.
Finally, for the complexity of Algorithm 3, the first two terms correspond to the cost of line 3. The computation of the canonical vectors ${\bf c}_{k}^{(\ell)}$ in line 4 requires $\mathcal{O}(m\log m)$ operations by using FFT, explaining the third term in the estimate. The complexity of the RHOSVD truncation is the sum of the costs of the truncated SVD of the side matrices, amounting to $\mathcal{O}(mR\min\{m,R\})$, and the cost of reconstruction of the Tucker core tensor $\bm{\beta}$, which is $\mathcal{O}(Rr^{3})$, see [35] for more details.
We notice that there is the traditional trade-off between the accuracy and computational complexity of the numerical algorithms. With regard to this concern, we comment that given $\varepsilon>0$, the polynomial degree $m=m(\varepsilon)$ depends on the regularity of input function $f(x)$ and may increase dramatically for non-regular functions having, for example, multiple local singularities or cusps, see for example Table 1. This is the case for interaction potentials of many-particle systems typically arising in quantum computational chemistry and bio-molecular modeling. For such challenging problems, the successful application of our method is based on the use of range-separated (RS) tensor formats, which will be discussed in what follows.
4 Applications to multi-particle modeling
In this section, we apply our hybrid tensor format to the challenging problem of approximating multi-particle interaction potentials. We consider the total potential generated by an $N$-particle system in $\mathbb{R}^{3}$, given by a weighted sum
(33) $P(x)={\sum}_{v=1}^{N}z_{v}p(\|x-x_{v}\|),\quad z_{v}\in\mathbb{R}\text{ and }x_{v},x\in[-1,1]^{d},$
where $x_{v}$ is the position of the $v$-th particle, $z_{v}$ is its weight and $p(\|x\|)$ is a radially symmetric generating kernel. These kernels are typically slowly decaying and singular at the origin. Examples include the Newton (Coulomb) $1/\|x\|$, Slater $\mathrm{e}^{-\lambda\|x\|}$, and Yukawa $\mathrm{e}^{-\lambda\|x\|}/\|x\|$ potentials, as well as Matérn covariance functions arising in various applications. Such practically interesting kernels have local singularities or cusps at the particle centers, and at the same time possess a non-local behavior that is responsible for the global interaction between all the particles in the system. Due to this short-long range behavior of the total collective potential $P(x)$ it does not allow for a low-rank tensor approximation in traditional tensor formats with satisfactory accuracy. This difficulty can be resolved by applying the RS tensor format [6].
In what follows, we will focus on the ubiquitous Newton kernel $p(\|x\|)=1/\|x\|$ for $d=3$, arising in electronic structure calculations and in bio-molecular modeling. We first recall shortly the results on a sinc-quadrature based constructive CP tensor representation of a single Newton kernel discretized on a 3D Cartesian grid [28, 11], thereby providing the CP tensor input for our mesh-free ChebTuck approximation Algorithm 3. We then introduce the idea of range-separation for efficient numerical treatment of the single Newton kernel [6]. Finally, we discuss how the total potential Eq. 33 can be constructed from the CP tensor representation of the single Newton kernel, for both bio-molecular modeling and lattice-structured compounds with vacancies, for which cases our ChebTuck approximation can be gainfully applied.
4.1 ChebTuck approximation for Newton kernel
4.1.1 A low-rank CP tensor representation of the Newton kernel
We first recall the constructive low-rank representation of the Newton kernel based on sinc-quadrature, as developed in [28, 11], and refer to these works for the detailed error analysis and numerically optimized procedure for rank minimization.
Let ${\bf P}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be the projection-collocation tensor of the Newton kernel $p(x)=1/\|x\|$ discretized on a $n_{1}\times n_{2}\times n_{3}$ 3D uniform Cartesian grid in $1,1]^{3}$, where the entries are defined by
(34) ${\bf P}_{{\bf i}}=\int_{[-1,1]^{3}}\dfrac{\psi_{{\bf i}}(x)}{\|x\|}\operatorname{d}x=\dfrac{1}{h_{1}h_{2}h_{3}}\int_{t_{i_{1}-1}^{(1)}}^{t_{i_{1}}^{(1)}}\int_{t_{i_{2}-1}^{(2)}}^{t_{i_{2}}^{(2)}}\int_{t_{i_{3}-1}^{(3)}}^{t_{i_{3}}^{(3)}}\dfrac{1}{\|x\|}\operatorname{d}x_{1}\operatorname{d}x_{2}\operatorname{d}x_{3},$
with the multi-index ${\bf i}=(i_{1},i_{2},i_{3})\in\{1,\cdots,n\}^{3}$, where $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ are the grid points in each dimension and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size. The basis function $\psi_{{\bf i}}$ is a product of piecewise constant functions in each dimension $\psi_{{\bf i}}(x)=\psi_{i_{1}}(x_{1})\psi_{i_{2}}(x_{2})\psi_{i_{3}}(x_{3})$ where $\psi_{i_{\ell}}(x_{\ell})=\chi_{[t_{i_{\ell-1}}^{(\ell)},t_{i_{\ell}}^{(\ell)})}(x_{\ell})/h_{\ell}$. To ease notation, we assume that $n_{1}=n_{2}=n_{3}=n$ in the following, and therefore use $h$ and $t_{i_{\ell}}$ to denote the grid size and grid points in all dimensions respectively. A numerically optimized procedure for rank minimization of the sinc-based CP approximation of $\bf P$} is then applied, leading to the final rank-$R$ CP format ${\bf P}_{R}$ of the collocation tensor $\bf P$}
(35) ${\bf P}\approx{\bf P}_{R}={\sum}_{k=1}^{R}{\bf p}_{k}^{(1)}\otimes{\bf p}_{k}^{(2)}\otimes{\bf p}_{k}^{(3)}$
where ${\bf p}_{k}^{(\ell)}$ are canonical vectors. We remark here that the canonical vectors in each dimension are the same due to the radial symmetry of the Newton kernel, i.e., ${\bf p}_{k}^{(1)}={\bf p}_{k}^{(2)}={\bf p}_{k}^{(3)}={\bf p}_{k}$. Note that here, the rank $R$ is adapted to the chosen approximation error and to the corresponding grid size $n$, and scales only logarithmically in both parameters. Similarly, the CP tensor decomposition can be constructed for a wide class of analytic kernels as mentioned in [28, 11]. In Fig. 1 left, we plot the canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\ldots,R$ of the CP approximation ${\bf P}_{R}$ of $\bf P$} discretized on a $n\times n\times n$ Cartesian grid for $n=256$, presented for example in [37, 34]. In this case, $R=26$ gives a relative error in the order of $10^{-7}$. We can see that there are canonical vectors localized around the singularity at the origin as well as slowly decaying ones, which clearly corresponds to the long-range and short-range behavior of the Newton kernel. We remark here that the number of vectors $R$ depends logarithmically on the size of the $n\times n\times n$ 3D Cartesian grid and the parameter $\varepsilon>0$ controlling the accuracy of the sinc-approximation. For increasing grid size $n$, the number of vectors with localized support around the singularity at the origin becomes larger.
Figure 1: Left: the canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\ldots,R$, $R=26$, along the $x$-axis for the rank-$R$ CP tensor approximation of collocation tensor of size ${\bf P}_{R}\in\mathbb{R}^{n\times n\times n}$ with $n=256$. Right: canonical vectors ${\bf a}_{k}^{(1)}$ of the long-range part in the total potential ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$ with $N=500$ particles and grid size $n=256$ for $k=1,\ldots,8$.
This low-rank CP tensor representation of the Newton kernel has been successfully used by two of the coauthors in high accuracy electronic structure calculations [39, 34]. In particular, it was applied for tensor based computation of the Hartree potential (3D convolution of the electron density with the Newton kernel), the nuclear potential operators, and the two-electron integral tensor, see [34] and the references therein. All computations have been performed using huge 3D $n\times n\times n$ Cartesian grids with $n$ of the order of $10^{4}\sim 10^{5}$, providing accuracy close to the results from the analytically based quantum chemical packages.
4.1.2 Why Chebyshev interpolation fails to approximate ${\bf P}_{R}$ without range-separation
In this subsection, we consider the ChebTuck approximation Algorithm 3 applied to the CP tensor ${\bf P}_{R}$ of the singular Newton kernel and discuss why it fails to give satisfactory results, which necessitates the use of the range-separated (RS) tensor format [6].
First, we compute the ChebTuck approximation relative error of the CP tensor ${\bf P}_{R}$ at the middle slice for different grid size $n$ and Chebyshev degree $m$, i.e.,
(36) $\mathrm{err}:=\max_{i_{1},i_{2}}\left|{\bf P}_{R}(i_{1},i_{2},n/2)-\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})\right|/\max_{i_{1},i_{2}}|{\bf P}_{R}(i_{1},i_{2},n/2)|,$
as shown in Table 1. We can see that in this case, only when $m\gg n$ the ChebTuck approximation provides satisfactory results, which is not practical and does not reduce the storage cost.
$n$ $m$ 129 257 513 1025 2049 4097 8193 16385
256 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6 1.6e-6 3.5e-7
512 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6 1.6e-6
1024 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6
2048 1.06 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4
4096 1.16 1.06 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4
Table 1: Relative $\ell_{\infty}$ error in the middle slice of the ChebTuck approximation of the single Newton kernel as in Eq. 36 for different grid size $n$ and Chebyshev degree $m$.
To have a closer look at the behavior, we take the collocation tensor of a single Newton kernel ${\bf P}_{R}\in\mathbb{R}^{n\times n\times n}$ with grid size $n=8192$ as an example. In this case, we get a rank $R=39$ CP approximation with relative error $10^{-7}$. We compute univariate Chebyshev interpolants of all canonical vectors ${\bf p}_{k}$ as described in Section 3.2.2 of degree $m$, denoted by $g_{k}(x)$. For each canonical vector ${\bf p}_{k}$, we measure the Chebyshev approximation error at the uniform grid points, i.e.,
(37) $\mathrm{err}:=\max_{i}\left|{\bf p}_{k}(i)-g_{k}(t_{i})\right|/\max_{i}|{\bf p}_{k}(i)|,$
for different Chebyshev degree $m$. The results are shown in Fig. 2. We can see that canonical vectors representing the long-range behavior of the Newton kernel, i.e., ${\bf p}_{1},\ldots,{\bf p}_{15}$ can be well approximated by low-degree Chebyshev interpolants (left), while for short-range canonical vectors, the convergence is much slower (middle). This is as expected since the short-range canonical vectors have singularities at the origin and require high-degree polynomials to capture the behavior. In Fig. 2 right, we plot the Chebyshev approximation error of different canonical vectors for a fixed Chebyshev degree. We can see given a fixed degree $m=129$, the error of the short-range functions is much larger than for the long-range functions with a jump around $k=15$. The same behavior can also be seen from Fig. 3, in which the absolute values of the Chebyshev coefficients of the interpolant of a long-range vector ${\bf p}_{10}$ and a short-range vector ${\bf p}_{30}$ are shown. We can see that for the long-range vector, the Chebyshev coefficients decay exponentially while for the short-range vector, the coefficients do not decay very much.
This motivates us to perform range-separation of the CP tensor ${\bf P}_{R}$ and use the range-separated tensor format [6] first before applying the ChebTuck approximation, which will be discussed in the next subsection.
Figure 2: Relative $\ell_{\infty}$ error Eq. 37 of the Chebyshev interpolation of the canonical vectors vs. the degree of the Chebyshev interpolant. Each line represents the error of the Chebyshev interpolant of a canonical vector ${\bf p}_{k}$ for $k=1,\ldots,15$ (left) and $k=16,\ldots,39$ (middle). Right: Relative $\ell_{\infty}$ error of the degree $m=129$ Chebyshev interpolation of the canonical vectors vs. the index $k$ of the univariate function.
Figure 3: Absolute values of the Chebyshev coefficients of the degree-129 Chebyshev interpolant of ${\bf p}_{10}$ (left) and ${\bf p}_{30}$ (right). Only coefficients of even degree are shown, i.e., ${\bf c}_{k}(0),{\bf c}_{k}(2),\ldots,{\bf c}_{k}(128)$ since $p(\|x\|)$ is even and ${\bf c}_{k}(1),\ldots,{\bf c}_{k}(129)=0$
4.1.3 Range-separation of the CP tensor ${\bf P}_{R}$
For a given appropriately chosen range-separation parameter $R_{s}\in[1,R]\cap\mathbb{Z}$ (see [6] for details of the strategies for choosing $R_{s}$), we further use the range-separated representation of the CP tensor ${\bf P}_{R}$ constructed in [6] in the form ${\bf P}_{R}={\bf P}_{R_{s}}+{\bf P}_{R_{l}}$, where
(38) ${\bf P}_{R_{s}}={\sum}_{k=1}^{R_{s}}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k},\ {\bf P}_{R_{l}}={\sum}_{k=R_{s}+1}^{R}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k}.$
Here, Eq. 38 represents the sums over the sets of indexes for the long- and short-range canonical vectors in the range-separation splitting of the Newton kernel. Then the short-range part ${\bf P}_{R_{s}}$ is a highly localized tensor which has very small support and can thus be stored sparsely (see [6] for details), while the ChebTuck approximation can be gainfully applied to the long-range part ${\bf P}_{R_{l}}$.
We apply Algorithm 3 to the long-range part ${\bf P}_{R_{l}}$ and measure the errors of the approximation at the middle slice as in Eq. 36 for different grid size $n$ and Chebyshev degree $m$, which are shown in Table 2. As opposed to Table 1, we can see that the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ can give satisfactory results already for reasonable Chebyshev degrees $m$. We also show the optimal Tucker ranks, i.e., size of the core tensor $\bm{\beta}$ of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ computed by RHOSVD with a relative error of $10^{-7}$ for different grid size $n$ and Chebyshev degree $m$ in the last column of Table 2. For fixed Chebyshev degree, the Tucker rank of the long-range part only increases logarithmically with the grid size $n$. This indicates that the long-range part can be well approximated by a ChebTuck format with a small number of parameters.
As a side note, we remark that the numerical evidence in Sections 4.1.2 and 4.1.3 aligns with the intuition underlying the fast multipole method (FMM) [26]: the far-field (long-range) component of the potential admits an accurate approximation by low-degree polynomials.
$n$ $m$ 129 257 513 1025 2049 4097 Tucker rank
256 8.4e-8 7.4e-8 7.3e-8 4.9e-9 1.6e-9 1.8e-10 9
512 5.2e-8 5.2e-8 4.6e-8 4.6e-8 3.0e-9 9.9e-10 11
1024 9.4e-9 9.4e-9 9.4e-9 8.7e-9 8.6e-9 5.7e-10 12
2048 3.6e-9 1.7e-9 1.7e-9 1.7e-9 1.6e-9 1.6e-9 13
4096 1.1e-4 7.5e-10 7.4e-10 7.4e-10 7.4e-10 7.1e-10 15
Table 2: Relative $\ell_{\infty}$ error as in Eq. 36 and in the middle slice and the Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ for different grid sizes $n$ and degree $m$. The Tucker rank is the same for different $m$ for a fixed grid size $n$, therefore only one value is shown.
4.2 ChebTuck approximation for multi-particle potentials
In this section, we first recall the construction of the range-separated tensor format [6] for the total potential Eq. 33 in an $N$-particle system. Then we consider two practical examples, a bio-molecular system and a lattice-structured compound with vacancies, where we apply the ChebTuck approximation to the long-range part of the total potential.
The total potential of an $N$-particle system $P(x)$ in Eq. 33 is nothing but a weighted sum of shifted Newton kernels. It was proven in [33] that the tensor representation of the collocation tensor of the total potential can also be obtained by shifting the collocation tensor of the Newton kernel, albeit through a more elaborate construction. In fact, it can be obtained through the application of shifting and windowing operators to the reference potential in Eq. 35, see [33], ${\cal W}_{\nu}={\cal W}_{\nu}^{(1)}\otimes{\cal W}_{\nu}^{(2)}\otimes{\cal W}_{\nu}^{(3)}$, $\nu=1,\cdots,N$, which applies to every particle located at $x_{\nu}$ in an $N$-particle system. Numerically it is obtained by construction of the single generating (reference) potential $\widetilde{\bf P}_{R}\in\mathbb{R}^{2n\times 2n\times 2n}$ in a twice larger computational box and shifting it according to the $x_{1},x_{2},x_{3}$ coordinates of every particle with further restricting (windowing) it to the original computational domain. Then the total multi-particle potential in Eq. 33 is computed as a projected weighted sum of shifted and windowed potentials ${\bf P}_{0}=\sum_{\nu=1}^{N}{z_{\nu}}\,{\cal W}_{\nu}(\widetilde{\bf P}_{R})\in\mathbb{R}^{n\times n\times n}$, see also [34] for more details.
It was proven in [6] that a tensor representation of this type in case of a multi-particle system of general type has almost irreducible large rank, $R_{N}=O(NR)$. To overcome this drawback, in [6] a new RS tensor format was introduced based on separation of the long- and short-range skeleton vectors in the generating tensor $\widetilde{\bf P}_{R}=\widetilde{\bf P}_{R_{s}}+\widetilde{\bf P}_{R_{l}}$ similar to the single Newton kernel version as in Section 4.1.3 which leads to a separable representation for the total sum of potentials,
${\bf P}_{0}={\sum}_{\nu=1}^{N}{z_{\nu}}\,{\cal W}_{\nu}(\widetilde{\mathbf{P}}_{R_{s}}+\widetilde{\mathbf{P}}_{R_{l}})=:{\bf P}_{s}+{\bf P}_{l}.$
Thanks to the RS tensor format, the CP-rank of the part of the sum, $\mathbf{P}_{l}$, containing only the long-range components of the canonical vectors of the shifted kernels, scales as $R_{l}\in O(R\log N)$, i.e., logarithmically in the number of particles in a molecular system [6]. Such rank compression of the long-range part $\mathbf{P}_{l}$ is achieved by applying the RHOSVD method for the canonical to Tucker (C2T) transform, and the subsequent Tucker to canonical (T2C) decomposition [39, 38]. This yields the resulting canonical tensor $\mathbf{P}_{R_{l}}$, ${\mathbf{P}}_{l}\mapsto{\mathbf{P}}_{R_{l}}$ with the reduced rank $RN\mapsto R_{l}\in O(R\log N)$.
One can aggregate the sum of short-range tensors ${\bf P}_{s}$ (having local supports which do not overlap each other) into the so called “cumulated canonical tensors”, $\mathbf{P}_{s}$, parametrized only by the coordinates of particle centers and parameters of the localized reference tensor, $\widetilde{\mathbf{P}}_{R_{s}}$. Above considerations resulted in the following definition (see [6] for more details).
Definition 4.1.
(RS-canonical tensors [6]). Given a reference tensor ${\bf A}_{0}$ supported by a small box such that $\mbox{rank}({\bf A}_{0})\leq R_{0}$, the separation parameter $\gamma\in\mathbb{N}$, and a set of distinct points ${\bf x}_{\nu}\in\mathbb{R}^{d}$, $\nu=1,\ldots,N$, the RS-canonical tensor format specifies the class of $d$-tensors ${\bf A}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, which can be represented as a sum of a rank-${R}_{l}$ canonical tensor ${\bf A}_{R_{l}}$ and a cumulated canonical tensor $\widehat{\bf A}_{S}$,
(39) ${\bf A}_{R_{l}}={\sum}_{k=1}^{R_{L}}\xi_{k}{\bf a}_{k}^{(1)}\otimes\cdots\otimes{\bf a}_{k}^{(d)},\ \widehat{\bf A}_{S}={\sum}_{\nu=1}^{N}c_{\nu}{\bf A}_{\nu}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}},$
where $\widehat{\bf A}_{S}$ is generated by replication of the reference tensor ${\bf A}_{0}$ to the points ${\bf x}_{\nu}$. Then the RS canonical tensor is represented in the form ${\bf A}={\bf A}_{R_{l}}+\widehat{\bf A}_{S}$, where $\mbox{diam}(\mbox{supp}{\bf A}_{\nu})\leq 2\gamma$ in the index size.
In what follows, we consider biomolecular systems corresponding to truncated fragments of the protein Fasciculin 1 with different numbers of atoms $N$. Fasciculin 1 is an anti-acetylcholinesterase toxin from the green mamba snake venom comprising 1,228 atoms [5, 41]. Throughout this section and Sections 4.2.1 and 4.2.2, the term biomolecule refers to this specific system.
In Fig. 4, we visualize the middle slice for the long-range part of the RS tensor, see the left figure, representing the total electrostatic potential of the biomolecule with $N=500$ particles, i.e., ${\bf A}_{R_{l}}$ of the $N$-particle potential in CP tensor format, with the reduced canonical rank $R_{l}=506\sim R\log(N)$ as compared to the original canonical rank $R_{N}=7000\sim NR$ computed by the RHOSVD-based rank reduction method [39]. We also show the error of the canonical-to-Tucker transform and of the consequent Tucker-to-canonical transform, ultimately canonical-to-canonical transform, in the middle and right figures in Fig. 4, respectively. We observe a good approximation error recovering the accuracy of rank truncation in tensor transforms. In Figure 1 right, we also visualize the several canonical vectors ${\bf a}_{k}^{(1)}$, $k=1,\ldots,10$, in the $x$-direction of the long-range part ${\bf A}_{R_{l}}$ of the same potential as in Figure 4.
In the next sections, we demonstrate the similar accuracy control for the proposed mesh-free two-level tensor approximation applied to the sums of Newton kernels in the case of biomolecules as well as for the lattice type compounds.
Figure 4: Left: Middle slice ${\bf A}_{R_{l}}(:,:,n/2)$ of the (compressed) long-range part of the RS tensor for the electrostatic potential of the biomolecule with $N=500$ particles computed on the 3D $n\times n\times n$ Cartesian grid with $n=1024$ and rank $R_{l}=506$. Middle: the error of the canonical-to-Tucker transform, with $\varepsilon=10^{-6}$. Right: the total error of the canonical-to-canonical tensor transform with the $\varepsilon=10^{-5}$.
4.2.1 The biomolecule potential with $N=500$ particles
In this section, we consider the ChebTuck approximation to the long range part of the biomolecule with $N=500$ particles. In this case, the underlining function we want to approximate is given by a 3-dimensional canonical tensor ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$, which can be pre-computed for any given grid size $n$ as described in Section 4.2. In this case, Algorithm 1 is not applicable but Algorithms 2 and 3 can be used.
We first consider the runtime and accuracy of Algorithms 2 and 3 for different discretization grid sizes $n$. We set the Chebyshev degrees in each dimension to be $m_{1}=m_{2}=m_{3}=m=129$ and the Tucker-ALS (RHOSVD in case of Algorithm 3) truncation tolerance to be $10^{-6}$. We record the runtime and $\ell_{\infty}$ error at the middle slice for each case in Table 3. Note that since the trivariate cubic spline in Algorithm 2 is cubic in grid size $n$ and for each Chebyshev degree $m$ (see Proposition 3.17), we only compute the results for $n=256$. We observe that the accuracy of the algorithms is similar while Algorithm 3 is significantly more efficient as it leverages the CP tensor structure of the long-range part as well as the RHOSVD method for Tucker compression. This is consistent with the theoretical estimates in Proposition 3.17. The ChebTuck format also becomes more accurate with increasing grid size $n$ which provides more information for the underlying function to be approximated.
Algorithm 2 Algorithm 3
$\ell_{\infty}$ error runtime $\ell_{\infty}$ error runtime
$n=256$ $8.59\cdot 10^{-7}$ 9.44 $8.59\cdot 10^{-7}$ 0.07
$n=512$ $6.89\cdot 10^{-8}$ 0.09
$n=1024$ $4.21\cdot 10^{-9}$ 0.13
$n=2048$ $6.37\cdot 10^{-11}$ 0.18
Table 3: Runtime in seconds and $\ell_{\infty}$ error for the ChebTuck approximation of the potential of the biomolecule with $N=500$ particles in $\mathbb{R}^{3}$ for different grid sizes $n$.
Next, we visualize the ChebTuck approximation error as well as the RHOSVD compression error for the potential with $n=256$ as in Fig. 5. In the left figure, we show the middle slice of the approximated potential $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$. Note that here we abuse the notation and use $\hat{f}_{\bf m}(:,:,t_{n/2})$ to denote the matrix containing function values of $\hat{f}_{\bf m}$ at the uniform grid points in the middle slice, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})=[\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})]_{i_{1}i_{2}}$. In the middle and right figures, we show the error during RHOSVD compressing and the total ChebTuck approximation error, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$ and $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$, respectively. Note the difference between $\hat{f}_{\bf m}$ and $\tilde{f}_{\bf m}$ as described in Section 3.1. We observe good approximation accuracy of our ChebTuck format. Another notable observation is that the error during RHOSVD compression is much smaller than the prescribed tolerance $\varepsilon=10^{-7}$, which indicates that the singular values decay extremely fast after the first few significant ones. Similar figures are also obtained for larger grid sizes $n=512,1024,2048$ as shown in Fig. 6. Our ChebTuck format achieves even higher accuracy for larger grid sizes, consistent with the results in Table 3.
We also remark here that the choice of the univariate interpolations $q_{k}^{(\ell)}(x)$ in Eq. 19 is crucial for the accuracy of the ChebTuck approximation since the convergence rate of the Chebyshev interpolants $g_{k}^{(\ell)}(x)$ depends on the smoothness of the univariate interpolations $q_{k}^{(\ell)}(x)$, as discussed in Lemma 3.8. For example, in Fig. 7, we replace the spline interpolation in Eq. 19 by piecewise linear (middle) and nearest neighbor (right) interpolation and observe that the ChebTuck format accuracy deteriorates significantly even for a large grid size $n=2048$.
Figure 5: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format of the biomolecule with $N=500$ and $n=256$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
Figure 6: The degree $m=129$ ChebTuck approximation total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the biomolecule for grid sizes $n=512,1024,2048$ from left to right.
Figure 7: The total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck approximation of the biomolecule for grid size $n=2048$ with different univariate interpolations in Eq. 19. Left: cubic spline, middle: piecewise linear, right: nearest neighbor.
4.2.2 $O(\log N)$ scaling of the Tucker rank of the ChebTuck format
In this section, we demonstrate numerically that the Tucker rank of our mesh-free ChebTuck format of the long-range part of the total potential in the biomolecule scales logarithmically with the number of particles $N$, consistent with discussions in Section 3.2.3 as well as the result for grid-based range-separated format as in [6, Theorem 3.1]. We consider the biomolecule in $d=3$ dimension with $N$ particles discretized on Cartesian grid with grid size $n=2048$. For varying particle number $N=100,\ldots,600$ and Chebyshev degree $m=3,5,9,\ldots,1025$, we show the Tucker rank of the ChebTuck format with $\varepsilon=10^{-7}$ RHOSVD truncation tolerance in Table 4. We can see that the Tucker rank grows only mildly with respect to the number of particles $N$. To have a closer look at the actual behavior, we also plot the singular values of the mode-$2$ side matrices of the Chebyshev coefficient tensor $\bf C$} (which is in CP format by construction) for $m=129$ in Fig. 8. We can see that the singular values drop to the level of machine precision after the first few significant ones, which also explains the behavior of the middle plot in Fig. 5.
NN $m$ 9 17 33 65 129 257 513 1025
100 9 17 26 26 26 26 26 26
200 9 17 30 30 30 30 30 30
300 9 17 31 32 32 32 32 32
400 9 17 31 32 32 32 32 32
500 9 17 31 32 32 32 32 32
600 9 17 31 33 33 33 33 33
Table 4: Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck format with RHOSVD truncation tolerance $10^{-7}$ for the long-range part of the total potential in a bio-molecule with $n=2048$ for different particle numbers NN and Chebyshev degrees $m$.
Figure 8: Left: Singular values of the mode-2 side matrices of the Chebyshev coefficient tensor ${\bf C}\in\mathbb{R}^{m\times m\times m}$ with $m=129$ for the long-range part of the total potential in the biomolecule with $n=2048$. Right: the zoom-in from the 10th to the 40th singular values of the same plot.
4.2.3 A lattice-type compound with vacancies
Finally, we apply our hybrid tensor format ChebTuck to the problem of low-parametric approximation of electrostatic potentials of large lattice-type structures, see [33] for tensor-based approach to lattice summation of electrostatic potentials. Fast computation of lattice potentials was since longer an important topic of numerical analysis because of various applications in crystal modeling, in quantum chemistry computations, in polymer simulations and many other fields. Traditional numerical methods for lattice summation are based on classical fast Ewald summation method via transition of the problem to the Fourier space [23, 14, 31], or, alternatively, by using the fast multipole method [26].
Here we demonstrate our new approach in application to the long-range part of the many-particles lattice-structured electrostatic potential. In this case, the charge centers $x_{v}$ in Eq. 33 are located on a regular grid. For a lattice-type structure of size $24\times 24\times 4$ with $N=2304$ particles, the canonical representation of its potential is given by a CP tensor of size $n\times n\times n$ with $n=960$ and rank $R=50$. We apply Algorithm 3 to approximate the long-range part of the potential with Chebyshev degrees $m_{1}=m_{2}=m_{3}=m=129$ and Tucker-ALS truncation tolerance $10^{-7}$. The results are shown in Fig. 9. We observe that the error is of magnitude $10^{-6}$, which is consistent with the prescribed tolerance and the previous results on the bio-molecule example. The numerical results confirm the efficiency of the new mesh-less hybrid tensor decomposition on the example of complicated multi-particle potential arising in quantum chemistry, biomolecule modeling, etc. Indeed, it demonstrates the same approximation error as for the established grid-based tensor approximation method applied on relatively large grids, required for the good resolution of complicated multi-particle potentials.
Figure 9: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format for the long-range part of a lattice structure with $N=2304$ and $n=960$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
5 Conclusions
We introduce and analyze a mesh-free two-level hybrid Tucker tensor format for approximation of trivariate functions on a hypercube with both high and reduced regularity, which combines the tensor product Chebyshev polynomial interpolation in the global computational domain with the ALS-based Tucker decomposition of the Chebyshev coefficient tensor. For functions with multiple cusps (or even stronger singularities), our method applies to the long-range component of the target function represented in low-rank RS tensor format. The Tucker rank optimization via ALS-type rank-structured decomposition of the coefficient tensor leads to nearly optimal $\varepsilon$-rank Tucker decomposition of the initial function with merely the same rank parameter as for the alternative grid-based methods. We present numerous numerical tests demonstrating the efficiency of the proposed techniques on the examples of many-particle electrostatic potentials in biomolecular systems and in lattice-structured compounds. In case of input functions presented in the canonical (CP) tensor format, our method needs only 1D-Chebyshev interpolations of the 1D-canonical vectors.
Finally, we notice that the presented techniques can be applied to the Green functions for various elliptic operators and can be extended to higher dimensions. In the latter case, the combination with cross-based tensor decompositions in TT, QTT or H-Tucker formats is necessary to avoid the curse of dimensionality.
Reproducibility statement
The full source code and data to reproduce the numerical experiments is available at https://github.com/bonans/ChebTuck.